OpenStudy (anonymous):

solve for the variable, check the result 3/6= x/8

OpenStudy (anonymous):

x=4

jimthompson5910 (jim_thompson5910):

3/6=x/8 3*8=6x 24 = 6x 24/6 = x 4 = x x = 4

jimthompson5910 (jim_thompson5910):

Note: in the second step, I'm cross multiplying.

OpenStudy (anonymous):

u da man